OpenStudy (anonymous):

Please help! the sum of the digits of a two digit number is 8, twice the ten's digit plus three times the unit's digit is 22. Finding the number would be 44. is this true or false?

OpenStudy (anonymous):

So, lets do this: If the number is 44, the tens digit (the leftmost) is 4, and the ones digits is 4. So, first thing is: the sum of the digits of a two digit number is 8 So, 4+4 = 8. therefore this is true! Next, twice the ten's digit plus three times the unit's digit is 22 So, if we do: 4*2 + 4*3 = 8+12 = 20 So this false. Therefore it is false.